# Second- and third-order properties of multidimensional Langevin equations

### v9

@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. We explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^2] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9].:
@@
The $n=0$ term contributes $-\mathds{1}$ to $\langle\widehat{\mathbf{A}}(T)\rangle$. For the contribution of the $n=1$ term, we have from Gaussianity and Isserlis’s theorem:
@@
This gives:
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
In the statistical literature, the term “integrated variable” in a stochastic process refers to a variable with no stationary distribution, but whose increments, or increments of increments, etc., do possess a stationary distribution. Here, we focus on variables integrated of order one, i.e., variables whose increments possess a stationary distribution. Such variables can be used to model e.g. the position and orientation of a particle in a homogeneous, isotropic medium. The dynamics of these variables may depend on degrees of freedom represented by variables possessing a stationary distribution [^3] Our analysis does not explicitly address ratchet models where $\langle\dot{x}\mid x\rangle$ is periodic in $x$. Such cases can be handled by introducing a complex variable $w=e^{ix}$ if the period is taken to be $2\pi$.. For linear dynamics, this type of system is modeled as:
@@
where a possible additive constant in $\dot{y}$ has been omitted. In this case, the angular momentum is defined as:
@@
(64) $\dot{z}=\xi^{z},\quad\langle\xi^{z}(t)\rangle=0,\quad\langle\boldsymbol{\xi}^{\mathbf{x}}(t)\xi^{z}(t^{\prime})\rangle=2\mathbf{D}^{\mathbf{x}z}\delta(t-t^{\prime}),\quad\langle\xi^{z}(t)\xi^{z}(t^{\prime})\rangle=2D^{zz}\delta(t-t^{\prime}).$
The transformed diffusion coefficients are:
@@
If detailed balance is satisfied for $\mathbf{x}$, then $\mathbf{D}^{\mathbf{x}\mathbf{x}}(\mathbf{A}^{-1})^{\mathsf{T}}=-\mathbf{C}$ and $\mathbf{L}(\mathbf{x},z)=\mathbf{L}(\mathbf{x},y)$. Moreover, the detailed balance condition for $z$ simply becomes $\mathbf{D}^{\mathbf{x}z}=\mathbf{0}$.
@@
Note that in the case where $\mathbf{x}$ is one-dimensional, the stochastic rotation frequency always vanishes because of a zero relaxation rate, but the entropy production does not vanish if detailed balance is broken.
@@
Between two integrated variables, there is no additional quantity governing break of detailed balance. Thus the angular momentum between them is considered to be 0, assuming that dynamics do not depend in any way on the values of those variables—in particular, not any combination of them.
@@ 3.2. Quantitative significance of the deterministic contribution
The next question we wish to consider is that of quantitative significance. How do we judge whether or not the deterministic contribution $\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{x}$ to $y$ is significant? And how do we compare whether a fitted model is a good match to experimental data, at least to linear order? This question has been addressed in the case of the usual stationary variables: compare the values of the experimentally measured covariance functions to the theoretical model prediction, and if the difference is small compared to $\mathbf{C}$, it is declared a good fit. However, we cannot proceed in an analogous way because the measured quantities are differences $y(\tau)-y(0)$ whose expectations scale as $\tau^{1}$ whereas root-mean-square values $\langle(y(\tau)-y(0))^{2}\rangle^{1/2}$ scale as $\tau^{1/2}$. Hence, they are not comparable.
@@
$\dot{y}$ $=\alpha x+\xi^{y},$
where $\lambda>0$. The dimensionless combination involving $\alpha$ is $(\alpha/\lambda)\sqrt{D^{xx}/D^{yy}}$. We will explicitly arrive at this quantity by studying the function $\langle(y(\tau)-y(0))^{2}\rangle$. We introduce as before the variable $z$, with:
@@
For multidimensional $\mathbf{x}$, we may compare the square roots of the quadratic variation of the two terms in Eq. (63).
@@ 3.3. Comparison of experimental data with theoretical predictions
As an example, for a single (integrated) variable the quantity in Eq. (81) approaches as $\tau\to\infty$ the negative of the long-time diffusivity, to be compared to the short-time diffusivity [^4] It should be noted that the limit $\tau\to\infty$ cannot be taken for a trajectory of finite length. To illustrate this, suppose we have a trajectory of length $N+1$, and define $\Delta x_{n}:=x((n+1)\Delta t)-x(n\Delta t)$, $n=0,1,\ldots,N-1$, $\overline{\Delta x}:=\sum_{n=0}^{N-1}\Delta x_{n}/N$. We may estimate the velocity autocovariance function as $\widehat{C}(n)=\sum_{m=0}^{N-1-n}(\Delta x_{m}-\overline{\Delta x})(\Delta x_{m+n}-\overline{\Delta x})/N(\Delta t)^{2}$, $n=0,1,\ldots,N-1$ [16]. We might then estimate the long-time diffusivity by $(\Delta t)[\widehat{C}(0)/2+\sum_{n=1}^{N-1}\widehat{C}(n)]$. However, this is equal to zero. Thus, although the velocity autocovariance function is self-averaging for any given time lag, its integral is not. If we denote the upper limit of integration by $M$, then the limits $M\to\infty$, $N\to\infty$ do not commute..
@@ 4. Third-order properties
where we have suppressed $t$-dependence, and Einstein summation notation is used. Neither addition to the model can stand on its own: in general, $a^{i}_{jk}$ by itself leads to divergences of $\mathbf{x}\to\infty$, and $b^{ij}_{k}$ by itself renders $\langle\mathrm{d}[x^{i},x^{j}]/\mathrm{d}t\mid\mathbf{x}(t)\rangle$ not positive semidefinite. A third possible contribution,
@@
Thus $L(x^{i}x^{j},x^{k})$ gives complete information about probability currents to “third order” (counting two orders for quadratic $f(\mathbf{x})$, and one order for $x^{i}$).
@@
(90) $\widetilde{L}(x^{i},x^{j},x^{k})=\langle x^{i}(x^{j}\mathcal{A}^{k}(\mathbf{x})-\mathcal{A}^{j}(\mathbf{x})x^{k})\rangle+\frac{\langle\mathcal{B}^{ik}(\mathbf{x})x^{j}-\mathcal{B}^{ij}(\mathbf{x})x^{k}\rangle}{2}=\frac{L(x^{i}x^{j},x^{k})-L(x^{i}x^{k},x^{j})}{2}$
@@ 5. Inhomogeneous diffusion
The second task is to compute covariance functions. In the current model, $\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle=e^{\mathbf{A}\tau}\mathbf{x}(0)$, so:
@@
Now from $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$, we have:
@@
We can write down the differential equation in the other direction:
@@ 6.2. One-dimensional case
Now we introduce the stochastic Koopman operator $\mathcal{K}$ [17, 18]. It is defined as an operator on functions of state space $f(\mathbf{x})$, as follows:
@@
The significance of the eigenfunctions of the stochastic Koopman operator (hereafter called “Koopman eigenfunctions”) is that their dynamics follows a linear law in expectation. Techniques and results for estimation of the Koopman operator are presented in [19, 20]. When $a=b=0$, the Koopman eigenfunctions are the probabilist’s Hermite polynomials ${He}_{n}(x)$ with eigenvalues $-n$. In the present case, the Koopman eigenfunctions $f_{n}(x)$ and the corresponding eigenvalues $\lambda_{n}$ are:
@@
Now, just like the moments, $\langle{He}_{2}(x(\tau))x(0)\rangle=\mathcal{O}(a,b)$ and so
@@ 6.5. Integrated variables
In the case of an integrated variable $y$, where
(164) $\langle\dot{y}\mid\mathbf{x}\rangle=A^{y}_{i}x^{i}+a^{y}_{ij}(x^{i}x^{j}-C^{ij})$
@@ 6.8. Quantitative significance of third-order effects
Now, we address the problem of determining if the third-order quantities are quantitatively significant. Following the discussion for the second moments, we may consider the third moments to be significant when:
@@
Following the discussion of deviations of experiment from theory for the second moments, we may treat $(i,j,k)$ as a single index and similarly for $(i^{\prime},j^{\prime},k^{\prime})$, then multiply the l.h.s. by the matrix inverse of the r.h.s. The resulting matrix has a single non-zero eigenvalue $\mu$ equal to its trace. To satisfy symmetry, we may consider an ensemble of stochastic systems whose “covariance” of third moments (l.h.s. of the above) is:
@@
The $b$ coefficients may be considered significant when:
@@
As before, we may consider $(i,j)$ as a single index and similarly for $(i^{\prime},j^{\prime})$ and take the l.h.s. multiplied by the matrix inverse of the r.h.s. In this case, however, the resulting matrix will generally have multiple non-zero eigenvalues. Nevertheless, we may use the same arguments to deduce that the eigenvalues remain the same when symmetrizing the r.h.s. (which is inverted after neglecting permutations). In the absence of symmetry constraints, we may consider the “covariance” for an ensemble of stochastic systems to be:
@@
An alternative choice would be to divide the above by the dimension $d$. However, this choice is inconsistent with the scaling with the dimension of the “covariance” of the third moments. We now calculate the “significant” value for $b$ collectively:
(184) $\sqrt{b^{ij}_{k}b^{i^{\prime}j^{\prime}}_{k^{\prime}}D^{-1}_{ii^{\prime}}D^{-1}_{jj^{\prime}}C^{kk^{\prime}}}\sim\sqrt{\frac{d_{0}d(d+1)}{2}}.$
@@
where $(j,k)$ is considered as a single index, as is $(j^{\prime},k^{\prime})$. We need to invert the above matrix neglecting permutations of $(j,k)$ and $(j^{\prime},k^{\prime})$. This is done by introducing a matrix $N_{jkj^{\prime}k^{\prime}}$ symmetric under the swappings $j\leftrightarrow k$, $j^{\prime}\leftrightarrow k^{\prime}$, and $(j,k)\leftrightarrow(j^{\prime},k^{\prime})$ such that:
@@
We then have for the “significant” value:
(188) $\sqrt{a^{i}_{jk}a^{i^{\prime}}_{j^{\prime}k^{\prime}}M^{jkj^{\prime}k^{\prime}}(\mathbf{A}^{-1})^{l}_{i}(\mathbf{A}^{-1})^{l^{\prime}}_{i^{\prime}}C^{-1}_{ll^{\prime}}}\sim\sqrt{\frac{d_{0}^{2}(d_{0}+1)}{2}}.$
@@
We also need the “symmetric inverse” $\widetilde{N}$ of the above, neglecting permutations:
@@
This results in:
@@
Lastly, we may consider the angular momenta $L(x^{i}x^{j},x^{k})$. Motivated by the case of angular momenta between two variables, we may suppose a “covariance” of the form:
@@
where the expectation is evaluated in the linear model. It is seen that the above satisfies Eq. (85). Collectively, these quantities may be considered significant when:
(193) $\sqrt{L(x^{i}x^{j},x^{k})L(x^{i^{\prime}}x^{j^{\prime}},x^{k^{\prime}})C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}D^{-1}_{kk^{\prime}}}\sim 2\sqrt{(d_{0}+1)(d-1)D^{jj^{\prime}}C^{-1}_{jj^{\prime}}},$

### v11

@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) in [9] to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. We explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^7] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9].:
@@
The $n=0$ term contributes $-\mathds{1}$ to $\langle\widehat{\mathbf{A}}(T)\rangle$. For the contribution of the $n=1$ term, we have from Gaussianity and Isserlis’s theorem (also known as Wick’s theorem [16]):
@@
The remaining terms give higher powers of $T^{-1}$. This gives:
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
In the statistical literature, the term “integrated variable” in a stochastic process refers to a variable with no stationary distribution, but whose increments, or increments of increments, etc., do possess a stationary distribution. Here, we focus on variables integrated of order one, i.e., variables whose increments possess a stationary distribution. Such variables can be used to model e.g. the position and orientation of a particle in a homogeneous, isotropic medium. The dynamics of these variables may depend on degrees of freedom represented by variables possessing a stationary distribution [^8] Our analysis does not address ratchet models where $\langle\dot{x}\mid x\rangle$ is periodic in $x$. In the systems we consider, any dependence on a variable not possessing a stationary distribution is disallowed.. For linear dynamics, this type of system is modeled as:
@@
 [^9] As with $\mathbf{x}$, we have omitted a possible additive constant in $\dot{y}$. A system with an additive constant may be transformed into Eq. (59) by considering instead the variable $y(t)-ct$, where $c$ is a constant (in this case, $c=\langle\dot{y}\rangle$). This simplification is again used throughout. Note that such a constant must vanish when detailed balance is satisfied, since we must have in that case $\langle\dot{y}\rangle=0$.
where $\boldsymbol{\xi}^{\mathbf{x}}$ and $\xi^{y}$ are Gaussian random noise and all coefficients are constant. In this case, the angular momentum between $\mathbf{x}$ and $y$ is defined as:
@@
(66) $\dot{z}=\xi^{z},\quad\langle\xi^{z}(t)\rangle=0,\quad\langle\boldsymbol{\xi}^{\mathbf{x}}(t)\xi^{z}(t^{\prime})\rangle=2\mathbf{D}^{\mathbf{x}z}\delta(t-t^{\prime}),\quad\langle\xi^{z}(t)\xi^{z}(t^{\prime})\rangle=2D^{zz}\delta(t-t^{\prime}),$
where $\xi^{z}$ is Gaussian random noise, and the transformed diffusion coefficients are:
@@
If detailed balance is satisfied for $\mathbf{x}$ (i.e., $L(x^{i},x^{j})=0$), then $\mathbf{D}^{\mathbf{x}\mathbf{x}}(\mathbf{A}^{-1})^{\mathsf{T}}=-\mathbf{C}$ and $\mathbf{L}(\mathbf{x},z)=\mathbf{L}(\mathbf{x},y)$. Moreover, the detailed balance condition for $z$ simply becomes $\mathbf{D}^{\mathbf{x}z}=\mathbf{0}$.
@@
Note that in the case where $\mathbf{x}$ is one-dimensional, the stochastic rotation frequency always vanishes because of a zero relaxation rate, but the entropy production rate does not vanish if detailed balance is broken.
@@
Between two integrated variables, there is no additional quantity governing break of detailed balance. Thus the angular momentum between them is considered to be 0, assuming that dynamics do not depend in any way on the values of those variables.
@@ 3.2. Quantitative significance of the deterministic contribution
The next question we wish to consider is that of quantitative significance. How do we judge whether or not the deterministic contribution $\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{x}$ to $y$ is significant? And how do we compare whether a fitted model is a good match to experimental data, at least to linear order? This question has been addressed in the case of the usual stationary variables: compare the values of the experimentally measured covariance functions to the theoretical model prediction, and if the difference is small compared to $\mathbf{C}$, it is declared a good fit. However, we cannot proceed in an analogous way because the measured quantities are differences $y(\tau)-y(0)$ whose expectations scale as $\mathcal{O}(\tau^{1})$ whereas root-mean-square values $\langle(y(\tau)-y(0))^{2}\rangle^{1/2}$ scale as $\mathcal{O}(\tau^{1/2})$. Hence, they are not comparable.
@@
(74) $\dot{y}$ $=\alpha x+\xi^{y},$
where $\lambda>0$. The dimensionless combination involving $\alpha$ is $(\alpha/\lambda)\sqrt{D^{xx}/D^{yy}}$. We will explicitly arrive at this quantity by studying the function $\langle(y(\tau)-y(0))^{2}\rangle$. We introduce according to Eq. (65) the variable $z$, with:
@@
For multidimensional $\mathbf{x}$, we may compare the square roots of the quadratic variation of the two terms on the r.h.s. of Eq. (65).
@@ 3.3. Comparison of experimental data with theoretical predictions
As an example, for a single (integrated) variable the quantity in Eq. (84) approaches as $\tau\to\infty$ the negative of the long-time diffusivity, to be compared to the short-time diffusivity. It should be noted, however, that the limit $\tau\to\infty$ cannot be taken for a trajectory of finite length. To illustrate this, suppose we have an integrated variable $x$ sampled with time-step $\Delta t$, and define $\Delta x_{n}:=x((n+1)\Delta t)-x(n\Delta t)$ for integer $n$, and $\Delta x_{n}^{\prime}:=\Delta x_{n}-\langle\Delta x\rangle$ (the subscript $n$ in the expectation value is suppressed). The long-time diffusivity is:
(85) $\lim_{N\to\infty}\frac{\langle(x(N\Delta t)-x(0)-\langle x(N\Delta t)-x(0)\rangle)^{2}\rangle}{2N\Delta t}$ $=\frac{1}{2N\Delta t}\lim_{N\to\infty}\sum_{0\leq m<N}\sum_{0\leq n<N}\langle\Delta x_{m}^{\prime}\Delta x_{n}^{\prime}\rangle$
$=\frac{\Delta t}{2}\lim_{N\to\infty}\sum_{-N<k<N}\left({1-\frac{|k|}{N}}\right)C(k)$
$=\frac{\Delta t}{2}\sum_{k=-\infty}^{\infty}C(k)$
($m,n,k$ are integers), where $C(k):=\langle\Delta x_{n+k}^{\prime}\Delta x_{n}^{\prime}\rangle/(\Delta t)^{2}$ is the velocity autocovariance function. We may estimate the velocity autocovariance function from a trajectory of length $N+1$ by [17]:
(86) $\widehat{C}(k):=\frac{1}{N(\Delta t)^{2}}\sum_{\begin{subarray}{c}0\leq n<N\\ 0\leq n+k<N\end{subarray}}(\Delta x_{n+k}-\overline{\Delta x})(\Delta x_{n}-\overline{\Delta x}),$
where
(87) $\overline{\Delta x}:=\frac{1}{N}\sum_{0\leq n<N}\Delta x_{n}.$
However, substituting the estimate Eq. (86) into Eq. (85) gives:
(88) $\frac{\Delta t}{2}\sum_{-N<k<N}\widehat{C}(k)=\frac{1}{2N\Delta t}\sum_{0\leq m<N}\sum_{0\leq n<N}(\Delta x_{m}-\overline{\Delta x})(\Delta x_{n}-\overline{\Delta x})=0,$
where we re-indexed $m=n+k$. Thus, although the estimate Eq. (86) for the velocity autocovariance function is self-averaging for any given finite time lag, its sum over all time lags is not [^10] In the case where $\langle\Delta x\rangle$ is treated as known and can be used in place of $\overline{\Delta x}$ in Eq. (88), the resulting estimate of long-time diffusivity is $(N/2\Delta t)(\overline{\Delta x}-\langle\Delta x\rangle)^{2}$, which has the correct expectation value, but is not self-averaging.. If we denote the limits of summation by ±$\pm M$, then the limits $M\to\infty$, $N\to\infty$ do not commute.
@@ 4. Third-order properties
where $\mathbf{A}$, $\mathbf{D}$, $a^{i}_{jk}$, and $b^{ij}_{k}$ are constant coefficients, we have suppressed $t$-dependence on the r.h.s., and Einstein summation notation is used. Neither addition to the model can stand on its own: in general, $a^{i}_{jk}$ by itself leads to divergences of $\mathbf{x}\to\infty$, and $b^{ij}_{k}$ by itself renders $\langle\mathrm{d}[x^{i},x^{j}]/\mathrm{d}t\mid\mathbf{x}(t)\rangle$ not positive semidefinite. A third possible contribution,
@@
. Thus $L(x^{i}x^{j},x^{k})$ gives complete information about probability currents to “third order” (counting two orders for quadratic $f(\mathbf{x})$, and one order for $x^{i}$).
The relation Eq. (95) can also be used to prove Eq. (92) by using the vanishing of the divergence of the probability current:
(96) 00 0 $=-\int\mathrm{d}\mathbf{x}\,x^{i}x^{j}x^{k}\frac{\partial J^{l}(\mathbf{x})}{\partial x^{l}}$
$=\int\mathrm{d}\mathbf{x}\,J^{l}(\mathbf{x})\frac{\partial}{\partial x^{l}}(x^{i}x^{j}x^{k})$
$=\int\mathrm{d}\mathbf{x}\,\left({x^{i}x^{j}J^{k}(\mathbf{x})+x^{j}x^{k}J^{i}(\mathbf{x})+x^{k}x^{i}J^{j}(\mathbf{x})}\right).$
@@
(98) $\widetilde{L}(x^{i},x^{j},x^{k})=\langle x^{i}(x^{j}\mathcal{A}^{k}(\mathbf{x})-\mathcal{A}^{j}(\mathbf{x})x^{k})\rangle+\frac{\langle\mathcal{B}^{ik}(\mathbf{x})x^{j}-\mathcal{B}^{ij}(\mathbf{x})x^{k}\rangle}{2}=\frac{L(x^{i}x^{j},x^{k})-L(x^{i}x^{k},x^{j})}{2}.$
@@ 5. Inhomogeneous diffusion
The second task is to compute covariance functions. In the current model, $\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle=e^{\mathbf{A}\tau}\mathbf{x}(0)$, so by the law of iterated expectations:
@@
From $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$, we have:
@@
We can write down the differential equation for the covariance function in the other direction:
@@ 6.2. One-dimensional case
Now we introduce the (stochastic) Koopman operator $\mathcal{K}$ [18, 19]. It is defined as an operator on functions of state space $f(\mathbf{x})$, as follows:
@@
Although the dynamics of $\mathbf{x}$ may have arbitrary non-linearities, the Koopman operator is a linear operator. The significance of the eigenfunctions of the Koopman operator (hereafter called “Koopman eigenfunctions”) is that their dynamics follows a linear law in expectation. (Techniques and results for estimation of the Koopman operator are presented in [20, 21].) The Koopman operator $\mathcal{K}$ has the same eigenvalues as the Perron–Frobenius operator $\mathcal{P}$ appearing in the Fokker–Planck equation ($\partial p(\mathbf{x},t)/\partial t=\mathcal{P}p(\mathbf{x},t)$, where $\mathcal{P}$ operates on $\mathbf{x}$), because $\mathcal{K}$ is the adjoint of $\mathcal{P}$ [19] with respect to the inner product defined by:
(144) $\langle f,g\rangle:=\int\mathrm{d}\mathbf{x}\,f(\mathbf{x})^{*}g(\mathbf{x})$
(asterisk denoting the complex conjugate), which can be easily shown using Itô’s lemma (with suitable assumptions on $f,g$). When $a=b=0$, the Koopman eigenfunctions are the probabilist’s Hermite polynomials ${He}_{n}(x)$ with eigenvalues $-n$. In the present case, the Koopman eigenfunctions $f_{n}(x)$ and the corresponding eigenvalues $\lambda_{n}$ are:
@@
Now, just like the moments, $\langle{He}_{2}(x(\tau))x(0)\rangle=\mathcal{O}(a,b)$ [^13] This is easily seen by taking $a,b\to 0$ in which case $\langle{He}_{2}(x(\tau))x(0)\rangle\to e^{-2\tau}\langle{He}_{2}(x)x\rangle=0$. This quantity is odd in $x$, leading to the estimate $\mathcal{O}(a,b)$. An even quantity such as $\langle{He}_{3}(x(\tau))x(0)\rangle\to e^{-3\tau}\langle{He}_{3}(x)x\rangle=0$ would be $\mathcal{O}((a,b)^{2})$, by orthogonality of the Hermite polynomials. and so
@@ 6.5. Integrated variables
We now address the case of an integrated variable $y$, where
(173) $\langle\dot{y}\mid\mathbf{x}\rangle=A^{y}_{i}x^{i}+a^{y}_{ij}(x^{i}x^{j}-C^{ij}).$
@@ 6.8. Quantitative significance of third-order effects
Now, we address the problem of determining if the third-order quantities are quantitatively significant. Following the discussion for the second moments, we may consider a putative relation for the “ensemble covariance” of the third moments as:
@@
Following the discussion of deviations of experiment from theory for the second moments, we may treat $(i,j,k)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime},k^{\prime})$, then multiply the l.h.s. by the matrix inverse of the r.h.s. For the same reason as before, the resulting matrix has a single non-zero eigenvalue $\mu$ equal to its trace. To satisfy symmetry, we may consider an ensemble of stochastic systems whose “ensemble covariance” of third moments (l.h.s. of the above) is:
@@
A putative relation for the “ensemble covariance” of the $b$ coefficients can be written as:
@@
Again, we may consider $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$ and take the l.h.s. multiplied by the matrix inverse of the r.h.s. In this case, however, the resulting matrix will generally have multiple non-zero eigenvalues. Nevertheless, we may use the same arguments to deduce that the eigenvalues remain the same when symmetrizing the r.h.s. (which is inverted after neglecting permutations). In the absence of symmetry constraints, we may then consider the “ensemble covariance” for an ensemble of stochastic systems to be:
@@
An alternative choice would be to divide the above by the dimension $d$. However, this choice is inconsistent with the scaling with the dimension of the “ensemble covariance” of the third moments. We now calculate the “ensemble covariance” for $b$ collectively:
(193) $b^{ij}_{k}b^{i^{\prime}j^{\prime}}_{k^{\prime}}D^{-1}_{ii^{\prime}}D^{-1}_{jj^{\prime}}C^{kk^{\prime}}\sim\frac{d_{0}d(d+1)}{2}.$
@@
where $(j,k)$ is considered as a single “super-index”, as is $(j^{\prime},k^{\prime})$. We need to invert the above matrix neglecting permutations of $(j,k)$ and $(j^{\prime},k^{\prime})$. This is done by introducing a matrix $N_{jkj^{\prime}k^{\prime}}$ symmetric under the swappings $j\leftrightarrow k$, $j^{\prime}\leftrightarrow k^{\prime}$, and $(j,k)\leftrightarrow(j^{\prime},k^{\prime})$ such that:
@@
We then have the “ensemble covariance” for the $a$ coefficients collectively:
(197) $a^{i}_{jk}a^{i^{\prime}}_{j^{\prime}k^{\prime}}M^{jkj^{\prime}k^{\prime}}(\mathbf{A}^{-1})^{l}_{i}(\mathbf{A}^{-1})^{l^{\prime}}_{i^{\prime}}C^{-1}_{ll^{\prime}}\sim\frac{d_{0}^{2}(d_{0}+1)}{2}.$
@@
Similarly to before, we need the “symmetric inverse” $\widetilde{N}$ of the above, neglecting permutations:
@@
This results in the following “ensemble covariance”:
@@
Lastly, we may consider the angular momenta $L(x^{i}x^{j},x^{k})$. Motivated by the case of angular momenta between two variables, we may suppose an “ensemble covariance” of the form:
@@
where the expectation is evaluated in the linear model. It is seen that the above satisfies Eq. (92). We then have the “ensemble covariance” of these quantities collectively:
(202) $L(x^{i}x^{j},x^{k})L(x^{i^{\prime}}x^{j^{\prime}},x^{k^{\prime}})C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}D^{-1}_{kk^{\prime}}\sim 4(d_{0}+1)(d-1)D^{jj^{\prime}}C^{-1}_{jj^{\prime}},$